# How AI Ideas Affect the Creativity, Diversity, and Evolution of Human Ideas: Evidence From a Large, Dynamic Experiment

### v1

@@ 3.1.1. Choosing AUT Items
We had to choose a selection of items that people would brainstorm creative uses for. We chose five items for which the creativity classifier that we used had the highest accuracy. Previously, Organisciak et al. (2022) fine-tuned GPT-3 Davinci to predict the creativity of AUT items. This dataset contains 20,121 responses from 2,025 participants, across 21 distinct AUT items and nine distinct studies (Organisciak et al., 2022). [^6] We obtained this dataset by direct correspondence with Dr. Organisciak on February 23, 2023; the code that Dr. Organisciak used to generate this dataset is available at https://github.com/massivetexts/llm_aut_study/blob/main/notebooks/Process_AUT_GT.ipynb Each response was graded for creativity by humans and normalized to a scale of 1-5. Then Organisciak et al. (2022) fine-tuned GPT-3 Davinci on this dataset to predict creativity scores. Here, fine-tuning involves providing {Input (an AUT response), Output (human rating)} pairs to a pre-trained LLM. Then the LLM adjusts its parameters to produce a similar output given an input, proxying human judgements. Overall, the fine-tuned GPT-3 classifier had a correlation [^7] We obtained scores for this classifier by downloading the zip file from (https://github.com/massivetexts/llm_aut_study/blob/main/results/evaluation.zip), then navigating to gt_main2/gpt-ft-davinci-1.csv of $r=0.81$ with human judgment. Accuracy varied by item (Appendix Table 5). For our experiment, we picked the five items for which the classifier had the highest accuracy: tire (r=0.91), pants (r=0.91), shoe (r=0.91), table (r=0.9), and bottle (r=0.88). Our task is ‘in-domain’ for the classifier since we ask participants to do the same task for the same items the classifier was trained on.
@@ 3.1.2. Generating GPT Ideas
Our specific zero-shot prompt was informed by prior work on LLMs and creativity. Stevenson et al. (2022) administered the AUT to GPT-3 through a zero-shot prompt. However, this prompt generated much wordier responses $(M=24.1,SD=8.8)$ than the human responses in the Organisciak Dataset $(M=4.6,SD=5)$. Such a discrepancy would alert participants to what was AI vs. human generated, which would nullify the disclosure factor (whether the source of an idea is disclosed). Hence, we appended a request (Figure 2) to use roughly the same number of words (5) as the average human response.
Word Length Prompt Experiment
We conducted a Monte Carlo experiment to confirm our modified zero-shot prompt resulted in responses with a similar word length to humans. Using parameters from Stevenson et al. (2022)’s experiments: For $n=1000$ trials, we fixed presence penalty and frequency penalty at 1, randomly chose a temperature (higher values lead to more randomness) in [0.65, 0.7, 0.75, 0.8], and randomly chose one of our 5 AUT items. For each trial, ChatGPT generated five ideas. The modified prompt resulted in responses with an average word length (M=4.44, SD = 1.34) much closer to human responses (M=4.56, SD = 4.97) than the original zero-shot prompt (M = 25.38, SD = 8.55). A permutation test (Appendix B), further shows that this difference was significant at $p<0.001$. We used the ideas generated by our modified zero-shot prompt as stimuli for the main experiment.
@@ 4.1. Summary
Condition Number of AI Examples Number of Human Examples Source of Ideas Disclosed
Control
0 6 N
High Exposure;
Disclosed
4 2 Y
High Exposure;
Not Disclosed
4 2 N
Low Exposure;
Disclosed
2 4 Y
Low Exposure;
Not Disclosed
2 4 N
@@ 4.2. Ethics
The experiment was approved by the [REDACTED FOR REVIEW] IRB. The experiment was deemed exempt from ongoing oversight under Exemption 3: Benign Behavioral Intervention.
@@ 4.3. Recruitment
We recruited volunteer participants through three sources: (1) Facebook ads, (2) Reddit, and (3) the weekly newsletter of Creative Mornings [^8] https://creativemornings.com/ [^9] The first author contacted Creative Mornings, who agreed to include the experiment in the newsletter., which is ‘the world’s largest face-to-face creative community’. Creative Mornings is an organization geared towards creative professionals that organizes (e.g.) talks and meetups. We ensured all participants were above 18. While we did not offer monetary compensation, we offered to give participants information about themselves, such as their creativity relative to both humans and AI, and their ability to spot creative ideas. Providing information to participants about themselves is often effective for recruiting volunteer participants since it makes the task intrinsically rewarding (Reinecke and Gajos, 2015). Appendix F describes the information we provided to participants.
@@ 4.4.4. Task Instructions
See Appendix M for screenshots. We asked participants to rank ideas to ensure that they would engage with the example ideas. [^10] We did not use these rankings as a DV since—because a participant ranks the examples they are shown, and all examples are from the same condition—these ranks could not speak to between-condition differences, which is the focus of the paper. Depending on the condition, (1) either there were or were not AI ideas in this example set (exposure); (2) AI ideas were or were not labeled (disclosure). We use the same prompt for humans (the text under Submit Your Idea) as with ChatGPT (Stevenson et al., 2022), but with a slight modification to request a single idea. This prompt contains language consistent with best practices in divergent thinking assessment Beaty et al. (2021). After submitting an idea, participants received feedback on their idea’s uniqueness and how accurately they ranked the example ideas (Appendix M).
@@ 4.4.5. Response Chains
(Note that because human ideas are propagated at the {[condition], item} level, the human ideas in the [Control] condition are ‘clean’ from AI contamination. They were brainstormed after seeing sets of human-only ideas, also from the [Control] condition.)
@@ 6. Outcome Measures
We have three outcome measures (idea diversity, creativity, and AI adoption) and three levels of analysis (local, evolution, and global). See Table 4. The local level measures outcomes at the level of an individual trial (e.g., how a submitted response relates to example responses). The evolution level measures the rate of change of outcome variables with respect to the trial number in the response chain (i.e., experiment iteration). The global level compares all submitted responses in a condition to each other. For all pairwise comparisons, we use a Holm-Bonferroni adjustment for multiple comparisons. For idea diversity and AI adoption, we scale the dependent variable (cosine distance or cosine similarity, respectively) by 100 for easier interpretation.
@@
Creativity
How creative is the submitted response? This is measured by the prediction of the classifier from Organisciak et al. (2022).
Not applicable
Does the creativity of submitted ideas change over time? This is measured by the slope of the response chain’s trial number (i.e., iteration in the response chain) on creativity (the metric from Organisciak et al. (2022)).
Idea Diversity
How different is a participant’s response from example responses? This is measured by the median pairwise semantic distance between ideas a participant sees and their response.
How diverse were all the participant’s ideas in a condition? This is measured by the median pairwise distance between all submitted ideas in a condition.
Do ideas become more different from each other as the experiment goes on? We first measure the median pairwise distance (‘idea diversity’) of ideas at each trial number (i.e., iteration in the response chain). We then measure the slope of the trial number on idea diversity.
AI Adoption
How similar is a participant’s response to AI example responses? This is measured by the maximum semantic distance between a participant’s response and AI examples.
Not applicable
Not applicable
@@ 7.2.1. Local Level
Intuitively, local idea diversity is how different a response is from the examples a participant sees. There was no main effect of condition $(F(4,19.95)=0.09,p=0.98)$, and the effect of self-perceived creativity did not differ by condition $(F(4,2650.44)=1.59,p=0.18)$. But the effect of belief in AI’s relative creativity did differ by condition, $F(4,2635.80)=2.93,p=0.02$. As robustness checks, we ran the same specification with two alternative measures of idea diversity, mean pairwise distance and distance from the centroid. Regression results are broadly similar. See Appendix I.2 for regression results and pairwise comparisons.
We found mixed evidence that belief in AI’s relative creativity moderates local idea diversity. Regression results showed a small but significant interaction effect between the [High Exposure, Undisclosed] condition and relative AI creativity ($\beta=0.038$, $t(3244)=2.149$, 95% CI = $[0.003,0.073]$, $p=0.03$). We probed this effect with estimated marginal means, predicting local idea diversity for the bottom and top decile of participants by perception of AI creativity. Top-decile participants had slightly higher local idea diversity than bottom-decile participants in the [High Exposure, Undisclosed] condition ($\Delta=2.62,d=-0.34$) but although this difference was significant before multiple comparisons ($p=0.01$), it was not significant after adjusting for multiple comparisons, ($p=0.06$; see Appendix Table 17). Hence, we conclude there is mixed evidence for the role of belief in AI’s relative creativity as a moderator of local idea diversity.
@@ 7.3.1. Local Level
Exposure to AI ideas increased adoption for (self-perceived) high-creativity participants regardless of disclosure, but this was not the case for (self-perceived) low-creativity participants. There was a significant interaction between self-perceived human creativity and the [High Exposure, Disclosed] condition ($\beta=0.11$, $t(2588)=3.93$, 95% CI = $[0.06,0.17],p=0.0001$; Appendix Table 23). To probe this interaction, we used our model to predict AI adoption by condition for both the top 10% and bottom 10% of participants by self-perceived creativity (Figures 9, 9 and 9). For high-creativity participants (Figure 9), adoption rates appear to differ only by exposure (color) and not disclosure (shape). More formally, we tested whether the effect of exposure on adoption is larger when AI ideas are disclosed vs undisclosed. We find that for high-creativity participants, there is no difference in adoption between ([High Exposure, Undisclosed] - [Low Exposure, Undisclosed]) and ([High Exposure, Disclosed] - [Low Exposure, Disclosed]), $\Delta=1.69,d=-0.14,p=0.59$. That is, the effect of exposure is not moderated by disclosure. But for low-creativity participants, the difference in adoption for the undisclosed conditions ([High Exposure, Undisclosed] - [Low Exposure, Undisclosed]) was larger than the equivalent difference in adoption for disclosed conditions ([High Exposure, Disclosed] - [Low Exposure, Disclosed]), $\Delta=7.77,d=0.65,p=0.03$. That is, disclosing ideas as from AI reduced the effect of exposure on adoption for lower (self-reported) creativity participants. In summary, higher (self-reported) creativity people adopt AI ideas solely based on content, and not disclosure.
@@ 8. Discussion
On the other hand, the presence of AI ideas increased the diversity of human ideas. This is consistent with prior work suggesting collaborating with AI leads to more diverse or unexpected outputs (Yang et al., 2022; Osone et al., 2021; Lee et al., 2022; Branch et al., 2021; Gero and Chilton, 2019). Crucially, high AI exposure increased both average amounts of diversity and rates of change in idea diversity. The latter result is especially important. Small differences in rates of change can yield large aggregate differences over time. Future work—both simulations and dynamic experiments—can explore the implications of this increase in collective idea diversity unaccompanied by an increase in individual creativity. For instance, can this dynamic generate ‘innovation’?
Our finding around the evolution of diversity (Figure 6) is instructive. Seeing other people’s ideas reduced idea diversity in the control condition over time. This may suggest that successive participants were converging on a particular idea sequence. But then injecting AI ideas into the example set increased the diversity of submitted responses by partially ‘resetting’ this convergence. A promising avenue for future work: Can AI input reduce ‘groupthink’?
@@
8.3. Participants adopted AI ideas when the prompt was hard.
When AI ideas were labeled, participants were more likely to adopt AI ideas for difficult prompts rather than easy prompts. This finding is similar to what Roemmele (2021) observed, where seeing AI examples only influenced creative output when the task was difficult. Both our and Roemmele (2021)’s results are consistent with a more general finding that task difficulty is associated with increased reliance on automation (Goddard et al., 2014). However, since this finding is based on a small number of AUT prompts we view this finding as speculative.
If users turn to AI for difficult rather than trivial tasks, this would have several implications. On one hand, AI can augment human creativity where human imaginations falter. At the same time, do we really want our most difficult creative tasks to be assisted by LLMs? Separate from an ethical question is a practical one. Researchers raised concerns over ‘model collapse’ (Shumailov et al., 2023)—the deteriorating performance of LLMs when trained on their outputs. If reliance on AI for creative tasks becomes routine, this may contribute to model collapse, ironically decreasing the efficacy of such reliance. More work is needed to understand LLM reliance and task difficulty.
@@ 9. Limitations & Future Work
Our study has several limitations that can inform future work. First, we measured the effect of AI ideas for a single task. We chose this task because it is one of the most common creativity tasks (Abraham, 2016). But future work could explore if our results replicate for other kinds of tasks. Second, we had to operationalize ‘ChatGPT’ in some concrete way. The logic for our prompt was driven by ecological validity and prior work: We used a zero-shot prompt because that is what users would likely use, and the specific prompt we used was derived from prior research. We chose not to vary prompts in order not to further increase the complexity of an already complex experiment. Future work could explore if different prompts elicit different results. Another avenue for future work is only propagating the ‘best’ AI ideas forward. Third, future work should test if alternative classifiers or ways of conceiving variables yield different results. For idea diversity and AI adoption, we addressed this problem by showing that conceptually similar ways of measuring variables yielded qualitatively similar results. For our creativity measure, we used a highly accurate classifier (correlation with human judgments greater than $r=0.88$ for items we used) trained for this exact task, for these exact items. But of course, all models have some error and future work based on this model propagates these errors. Incidentally, human judges of creativity only correlate with other human judges at $r=0.88$ (Organisciak et al., 2023), suggesting the classifier we used may be approaching ‘the approximate ceiling at which we could expect a model to correlate with human judgments’ (Organisciak et al., 2023, pg. 11) of creativity. Fourth, our finding about AI adoption and task difficulty is based on five AUT items. Future work should explore this relationship with a larger number of stimuli. Fifth, we focus on one facet of creativity: originality. Future work can also explore whether AI ideas have different effects on other facets of creativity. Sixth, we employed a convenience sample of technology-interested users and creative professionals. While these two groups are most relevant to the phenomena in question, our sample also limits generalizability. Future work can explore these dynamics with different samples. Finally, we conducted this experiment close to the launch of ChatGPT. As AI becomes increasingly embedded in everyday life, attitudes towards AI and ways of engaging with AI may also change. Despite these limitations, our work offers the first large-scale, dynamic account of how ideas from LLMs affect collective thought.
10. Acknowledgements
We thank Joyce Chai, Misha Teplitskiy, and Harman Kaur for their feedback.

### v2

@@ 3.1.1. Choosing AUT Items
We had to choose a selection of items that people would brainstorm creative uses for. We chose five items for which the creativity classifier that we used had the highest accuracy. Previously, Organisciak et al. (2022) fine-tuned GPT-3 Davinci to predict the creativity of AUT items. This dataset contains 20,121 responses from 2,025 participants, across 21 distinct AUT items and nine distinct studies (Organisciak et al., 2022). [^5] We obtained this dataset by direct correspondence with Dr. Organisciak on February 23, 2023; the code that Dr. Organisciak used to generate this dataset is available at https://github.com/massivetexts/llm_aut_study/blob/main/notebooks/Process_AUT_GT.ipynb Each response was graded for creativity by humans and normalized to a scale of 1-5. Then Organisciak et al. (2022) fine-tuned GPT-3 Davinci on this dataset to predict creativity scores. Here, fine-tuning involves providing {Input (an AUT response), Output (human rating)} pairs to a pre-trained LLM. Then the LLM adjusts its parameters to produce a similar output given an input, proxying human judgments. Overall, the fine-tuned GPT-3 classifier had a correlation [^6] We obtained scores for this classifier by downloading the zip file from (https://github.com/massivetexts/llm_aut_study/blob/main/results/evaluation.zip), then navigating to gt_main2/gpt-ft-davinci-1.csv of $r=0.81$ with human judgment. Accuracy varied by item (Appendix Table LABEL:aut_desc_stats). For our experiment, we picked the five items for which the classifier had the highest accuracy: tire (r=0.91), pants (r=0.91), shoe (r=0.91), table (r=0.9), and bottle (r=0.88). Our task is ‘in-domain’ for the classifier since we ask participants to do the same task for the same items the classifier was trained on.
@@ 3.1.2. Generating GPT Ideas
Our specific zero-shot prompt was informed by prior work on LLMs and creativity. Stevenson et al. (2022) administered the AUT to GPT-3 through a zero-shot prompt. However, this prompt generated much wordier responses $(M=25.4,SD=8.5)$ than the human responses in the Organisciak Dataset $(M=4.6,SD=5)$. Such a discrepancy would alert participants to what was AI vs. human generated, which would nullify the disclosure factor (whether the source of an idea is disclosed). Hence, we appended a request (Figure 2) to use roughly the same number of words (5) as the average human response. The modified prompt resulted in responses with an average word length $(M=4.4,SD=1.3)$ much closer to human responses than the un-modified Stevenson prompt, $p<0.001$ by two-tailed permutation tests. We use the ideas from this modified Stevenson prompt for our experiments (Appendix B for more details).
@@ 4.1. Summary
Condition Number of AI Number of Human Source Disclosed
Examples Examples (Y/N)
Control 0 6 N
High Exposure; Disclosed 4 2 Y
High Exposure; Not Disclosed 4 2 N
Low Exposure; Disclosed 2 4 Y
Low Exposure; Not Disclosed 2 4 N
@@ 4.2. Ethics
The experiment was approved by our university’s institutional review board.
@@ 4.3. Recruitment
We recruited volunteer participants through three sources: (1) Facebook ads, (2) Reddit, and (3) the weekly newsletter of Creative Mornings [^7] https://creativemornings.com/ [^8] The first author contacted Creative Mornings, who agreed to include the experiment in the newsletter., which is ‘the world’s largest face-to-face creative community’. Creative Mornings is an organization geared towards creative professionals that organizes (e.g.) talks and meetups. We ensured all participants were above 18. While we did not offer monetary compensation, we offered to give participants information about themselves, such as their creativity relative to both humans and AI and their ability to spot creative ideas. Providing information to participants about themselves is often effective for recruiting volunteer participants since it makes the task intrinsically rewarding (Reinecke and Gajos, 2015). Appendix F describes the information we provided to participants.
@@ 4.4.4. Task Instructions
See Appendix M for screenshots. We asked participants to rank ideas to ensure that they would engage with the example ideas. [^9] We did not use these rankings as a DV since—because a participant ranks the examples they are shown, and all examples are from the same condition—these ranks could not speak to between-condition differences, which is the focus of the paper. Depending on the condition, (1) either there were or were not AI ideas in this example set (exposure); (2) AI ideas were or were not labeled (disclosure). We use the same prompt for humans (the text under Submit Your Idea) as with ChatGPT (Stevenson et al., 2022), but with a slight modification to request a single idea. This prompt contains language consistent with best practices in divergent thinking assessment (Beaty et al., 2021). After submitting an idea, participants received feedback on their idea’s uniqueness and how accurately they ranked the example ideas (Appendix M).
@@ 4.4.5. Response Chains
Note that because human ideas are propagated at the {[condition], item} level, the human ideas in the [Control] condition are ‘clean’ from AI contamination. They were brainstormed after seeing sets of human-only ideas, also from the [Control] condition.)
@@ 6. Outcome Measures
We have three outcome measures (idea diversity, creativity, and AI adoption) and three levels of analysis (local, evolution, and global). See Table LABEL:big_table. The local level measures outcomes at the level of an individual trial (e.g., how a submitted response relates to example responses). The evolution level measures the rate of change of outcome variables with respect to the trial number in the response chain (i.e., experiment iteration). The global level compares all submitted responses in a condition to each other. For all pairwise comparisons, we use a Holm-Bonferroni adjustment for multiple comparisons. For idea diversity and AI adoption, we scale the dependent variable (cosine distance or cosine similarity, respectively) by 100 for easier interpretation.
@@
Creativity How creative is the submitted response? This is measured by the prediction of the classifier from Organisciak et al. (2022). Not applicable Does the creativity of submitted ideas change over time? This is measured by the slope of the response chain’s trial number (i.e., iteration in the response chain) on creativity (the metric from Organisciak et al. (2022)).
Idea Diversity How different is a participant’s response from example responses? This is measured by the median pairwise semantic distance between ideas a participant sees and their response. How diverse were all the participant’s ideas in a condition? This is measured by the median pairwise distance between all submitted ideas in a condition. Do ideas become more different from each other as the experiment goes on? We first measure the median pairwise distance (‘idea diversity’) of ideas at each trial number (i.e., iteration in the response chain). We then measure the slope of the trial number on idea diversity.
AI Adoption How similar is a participant’s response to AI example responses? This is measured by the maximum semantic distance between a participant’s response and AI examples. Not applicable Not applicable
@@ 7.2.1. Local Level
Intuitively, local idea diversity is how different a response is from the examples a participant sees. There was no main effect of condition $(F(4,19.95)=0.09,p=0.98)$, and the effect of self-perceived creativity did not differ by condition $(F(4,2650.44)=1.59,p=0.18)$. But the effect of belief in AI’s relative creativity did differ by condition, $F(4,2635.80)=2.93,p=0.02$. As robustness checks, we ran the same specification with two alternative measures of idea diversity, mean pairwise distance and distance from the centroid. Regression results are broadly similar. However, post-hoc estimated marginal means showed non-significant contrasts, so we refrain from interpreting this finding. See Appendix I.2 for a more in-depth discussion, regression results, and pairwise comparisons.
@@ 7.3.1. Local Level
Exposure to AI ideas increased adoption for (self-perceived) high-creativity participants regardless of disclosure, but this was not the case for (self-perceived) low-creativity participants. There was a significant interaction between self-perceived human creativity and the [High Exposure, Disclosed] condition ($\beta=0.11$, $t(2588)=3.93$, 95% CI = $[0.06,0.17],p=0.0001$; Appendix Table 23). To probe this interaction, we used our model to predict AI adoption by condition for both the top 10% and bottom 10% of participants by self-perceived creativity (Figures 9, 9 and 9). For high-creativity participants (Figure 9), adoption rates appear to differ only by exposure (color) and not disclosure (shape). More formally, we tested whether the effect of exposure on adoption is larger when AI ideas are disclosed vs undisclosed. We find that for high-creativity participants, there is no difference in adoption between ([High Exposure, Undisclosed] - [Low Exposure, Undisclosed]) and ([High Exposure, Disclosed] - [Low Exposure, Disclosed]), $\Delta=-1.69,d=-0.14,p=0.59$. That is, the effect of exposure is not moderated by disclosure. But for low-creativity participants, the difference in adoption for the undisclosed conditions ([High Exposure, Undisclosed] - [Low Exposure, Undisclosed]) was larger than the equivalent difference in adoption for disclosed conditions ([High Exposure, Disclosed] - [Low Exposure, Disclosed]), $\Delta=7.77,d=0.65,p=0.03$. That is, disclosing ideas as from AI reduced the effect of exposure on adoption for lower (self-reported) creativity participants. In summary, higher (self-reported) creativity people adopt AI ideas solely based on content, and not disclosure.
@@ 8. Discussion
On the other hand, the presence of AI ideas increased the diversity of human ideas. This is consistent with work suggesting collaborating with AI leads to more diverse or unexpected outputs (Yang et al., 2022; Osone et al., 2021; Lee et al., 2022; Branch et al., 2021; Gero and Chilton, 2019) and inconsistent with other work that finds collaborating with LLMs decreases diversity (Padmakumar and He, 2024; Doshi and Hauser, 2024; Dell’Acqua et al., 2023). But we highlight that our study is testing passive exposure to AI ideas and not active engagement with AI ideas. Our setup—passive exposure to AI ideas, scattered amongst human ones—maps onto how many users now experience AI ideas. Hence, it may be that active engagement with LLMs decreases content diversity but simply seeing these ideas as ‘sparks’ (Gero et al., 2022) increases content diversity. And because many more people may be passively exposed to LLM outputs than actively engaging with LLMs, the effect of passive exposure is important to understand.
Crucially, high AI exposure increased both average amounts of diversity and rates of change in idea diversity. The latter result is especially important. Small differences in rates of change can yield large aggregate differences over time. Future work—both simulations and dynamic experiments—can explore the implications of this increase in collective idea diversity unaccompanied by an increase in average individual creativity. For instance, can this dynamic generate ‘innovation’?
Our finding around the evolution of diversity (Figure 6) is instructive. Seeing other people’s ideas reduced idea diversity in the control condition over time. This may suggest that successive participants were converging on a particular idea sequence. But then injecting AI ideas into the example set increased the diversity of submitted responses by partially ‘resetting’ this convergence. Our finding relates to recent work proposing AI systems that generate ‘alien’ scientific hypotheses humans would not think of (Sourati and Evans, 2023). More generally, a promising avenue for future work: Can AI input reduce ‘groupthink’?
@@
8.3. Participants may adopt AI ideas when the task is difficult
When AI ideas were labeled, participants were more likely to adopt AI ideas for difficult prompts rather than easy prompts. Although this finding should be taken as speculative (due to the small number of items), it is similar to what (Roemmele, 2021) observed, where seeing AI examples only influenced creative output when the task was difficult. Both our and (Roemmele, 2021)’s results are consistent with a theoretical account of task difficulty being associated with increased reliance on automation (Goddard et al., 2014). Future work can further test whether users adopt AI ideas for more challenging creative tasks.
If users turn to AI for difficult rather than trivial tasks, this has several implications. On one hand, AI can augment human creativity where human imaginations falter. At the same time, researchers raised concerns over ‘model collapse’ (Shumailov et al., 2023)—the deteriorating performance of LLMs when trained on their outputs. If reliance on AI for creative tasks becomes routine, this may contribute to model collapse, ironically decreasing the efficacy of such reliance.
@@ 9. Limitations & Future Work
Our study has several limitations that can inform future work. First, we measured the effect of AI ideas for a single task. We chose this task because it is one of the most common creativity tasks (Abraham, 2016). But future work could explore if our results replicate for other kinds of tasks. Second, we had to operationalize ‘ChatGPT’ in some concrete way. The logic for our prompt was driven by ecological validity and prior work: We used a zero-shot prompt because that is what users would likely use, and the specific prompt we used was derived from prior research. We chose not to vary prompts in order not to further increase the complexity of an already complex experiment. Future work could explore if different prompts elicit different results. Another avenue for future work is only propagating the ‘best’ AI ideas forward. Third, future work should test if alternative classifiers or ways of conceiving variables yield different results. For idea diversity and AI adoption, we addressed this problem by showing that conceptually similar ways of measuring variables yielded qualitatively similar results. For our creativity measure, we used a highly accurate classifier (correlation with human judgments greater than $r=0.88$ for items we used) trained for this exact task, for these exact items. But of course, all models have some error and future work based on this model propagates these errors. Incidentally, human judges of creativity only correlate with other human judges at $r=0.88$ (Organisciak et al., 2023), suggesting the classifier we used may be approaching ‘the approximate ceiling at which we could expect a model to correlate with human judgments’ (Organisciak et al., 2023, pg. 11) of creativity. Nonetheless, future work can adopt a similar design but with human judgments or different measures. Fourth, our finding about AI adoption and task difficulty is based on five AUT items. Future work should explore this relationship with a larger number of stimuli. Fifth, we focus on one facet of creativity: originality. Future work can also explore whether AI ideas have different effects on other facets of creativity. We also did not necessarily measure idea “quality”, which is distinct from originality and diversity. Sixth, we employed a convenience sample of technology-interested users and creative professionals. While these two groups are most relevant to the phenomena in question, our sample also limits generalizability. Future work can explore these dynamics with different samples. Seventh, these dynamics may differ for future LLMs. Finally, we conducted this experiment close to the launch of ChatGPT. As AI becomes increasingly embedded in everyday life, attitudes towards AI and ways of engaging with AI may also change. Despite these limitations, our work offers the first large-scale, dynamic account of how ideas from LLMs affect collective thought.